### v3

# Stealthy Patch-Wise Backdoor Attack in 3D Point Cloud via Curvature Awareness

Abstract
Backdoor attacks pose a severe threat to deep neural networks (DNNs) by implanting hidden backdoors that can be activated with predefined triggers to manipulate model behaviors maliciously. Existing 3D point cloud backdoor attacks primarily rely on sample-wise global modifications, which suffer from low imperceptibility. Although optimization can improve stealthiness, optimizing sample-wise triggers significantly increases computational cost. To address these limitations, we propose the Stealthy Patch-Wise Backdoor Attack (SPBA), the first patch-wise backdoor attack framework for 3D point clouds. Specifically, SPBA decomposes point clouds into local patches and employs a curvature-based imperceptibility score to guide trigger injection into visually less sensitive patches. By optimizing a unified patch-wise trigger that perturbs spectral features of selected patches, SPBA significantly enhances optimization efficiency while maintaining high stealthiness. Extensive experiments on ModelNet40 and ShapeNetPart further demonstrate that SPBA surpasses prior state-of-the-art backdoor attacks in both attack effectiveness and resistance to defense methods. The code is available at https://github.com/HazardFY/SPBA.
1 Introduction
Point clouds serve as a highly effective data structure for representing 3D geometric data [11], finding increasing utility in safety-critical applications including autonomous driving [17], healthcare [42], and robotics [49]. With this expanding deployment, the security of point cloud models becomes paramount, particularly concerning their vulnerability to backdoor attacks [16]. In these attacks, adversaries inject a pre-defined trigger into a small portion of training data. This manipulation causes the backdoored model, upon training, to consistently misclassify samples containing the trigger into a target class, while maintaining accurate predictions for benign inputs. A well-designed trigger remains inconspicuous within the data yet potently activates malicious behavior in the backdoored model. Therefore, trigger design is crucial for both the effectiveness and stealth of backdoor attacks [33, 45].
Figure 1: Comparison of latest sample-wise attacks and our stealthy patch-wise attack. Green points represent the original sample, while red points indicate modifications introduced by injected triggers.
Early explorations into 3D point cloud backdoor attacks utilized direct global geometric modifications as triggers, such as adding spherical point clusters or applying specific rotations [16, 37, 43]. These global modifications produce deformations readily discernible to the human eye and susceptible to standard data preprocessing techniques like Statistical Outlier Removal (SOR) [48] or rotation augmentation [7]. Recent efforts have shifted towards incorporating local context into trigger design to enhance stealthiness. For instance, IRBA [7] utilized a nonlinear and local transformation to generate sample-wise triggers, while IBAPC [5] employed global spectral triggers to balance global shape preservation with localized point perturbations. These triggers are more adaptable, fitting the local geometric structure of samples [7] while preserving the overall global geometry for enhanced stealthiness [5]. However, these methods do not sufficiently focus on local context, limiting their potential for improved stealthiness. Their triggers either alter the spatial relationship between local and global structures or induce broader deformations, even when aiming for local perturbations.
To this end, we aim to design a novel and highly imperceptible backdoor attack trigger that solely leverages local context information for subtle local modifications. In the 2D image domain, local triggers are typically designed as unified patch-wise triggers [10], which blend seamlessly into diverse image backgrounds. However, the absence of an inherent grid structure in 3D point clouds makes it challenging to design and inject consistent patch-wise deformation triggers without introducing noticeable anomalies or compromising geometric integrity. To address this challenge, we design the patch-wise trigger in the spectral domain rather than the raw spatial domain. Specifically, the Graph Fourier Transform (GFT) is employed to map point coordinates of local patches into the spectral domain, where the overall local geometric structure is maintained while the point-level fine-grained details are perturbed. As noted by [23], human perception varies across different regions of a 3D point cloud, with perturbations in geometrically complex areas being less perceptible. Building on this insight, we selectively filter out smooth patches and apply the patch-wise trigger exclusively to geometrically complex regions, thereby enhancing stealthiness, as illustrated in Fig. 1. Notably, the patch-wise spectral trigger requires significantly lower computational cost compared to the sample-wise spectral trigger in the previous IBAPC attack [5]. This greatly improves the efficiency and deployability of our attack method. What’s more, the enriched optimization landscape enabled by patch-wise decomposition, together with selected patch optimization, is empirically found to be beneficial for improving attack effectiveness.
In this paper, we propose the Stealthy Patch-Wise Backdoor Attack (SPBA), which introduces the first patch-wise trigger in 3D point cloud backdoor attack. Our approach first decomposes a point cloud sample into local patches, enabling finer control over trigger injection. To select geometrically complex patches, we design a curvature-based Patch Imperceptibility Score (PIS) to quantify patch complexity. The patches with high PIS are selected and then converted into KNN graphs, providing a structured representation of local geometric relationships. We employ the Graph Fourier Transform to map point coordinates into the spectral domain, where perturbations can be introduced in a controlled and structured manner. To maintain the geometric integrity of the affected patches, we apply an optimized spectral trigger that selectively modifies the spectral features while preserving the geometric structure. Finally, the Inverse Graph Fourier Transform (IGFT) is utilized to reconstruct the poisoned patches, which are then aggregated with the unmodified regions to form the final poisoned point cloud sample. An alternating optimization process is employed to jointly refine both the point cloud model and the patch-wise trigger, ensuring optimal effectiveness and stealthiness of the trigger. Our main contributions include:
  • We propose SPBA, the first patch-wise backdoor attack on 3D point clouds that leverages the Graph Fourier Transform (GFT) to enable efficient and precise spectral perturbations at the local level, while preserving the global geometric structure.
  • We introduce a curvature-based patch imperceptibility scoring mechanism to identify and select geometrically complex patches for trigger injection, enhancing the stealthiness and effectiveness of SPBA.
  • We conduct comprehensive experiments on two public benchmarks to demonstrate the superiority of SPBA over previous state-of-the-art backdoor attacks on 3D point clouds in terms of attack effectiveness, stealthiness, and efficiency.
2 Related Work
2.1 Patch-wise Backdoor Attacks in 2D Images
Different from adversarial attacks [4, 46], backdoor attacks pose a significant security threat in the training process by embedding hidden triggers during the training phase, enabling adversaries to manipulate model predictions during inference [20, 8]. The effectiveness of a backdoor attack heavily depends on trigger design, which can be broadly categorized into global triggers and local patch-wise triggers. Prior works have investigated global triggers from various perspectives, including image steganography [18, 32], frequency-domain injection [6], and optimization-based approaches [40], aiming to balance stealthiness and attack effectiveness. As for local patch-wise triggers, BadNets [10] first demonstrated the feasibility of backdoor attacks in the 2D image domain with a fix patch at the corner. Since then, patch-wise triggers have become a fundamental paradigm in backdoor attacks due to their high attack success rates and stealthiness. Chen  et al. [3] introduced a blending strategy, embedding triggers with partial transparency to make them less detectable. Yuan  et al. [43] found that Vision Transformers (ViTs) are particularly vulnerable to patch-wise triggers, due to their global dependencies within patches. Some works further explored the adaptability of patch-wise trigger to various training paradigms, including self-supervised learning [29] and multimodal contrastive learning [1]. Gao  et al. [9] further improved patch-wise attacks by designing a sparse and invisible trigger through bi-level optimization, minimizing visual detectability while preserving attack effectiveness. What’s more, Li  et al. [19] adopted patch-wise triggers as a harmless and stealthy way to protect the dataset copyright.
Intuitively, perturbing only a small subset of points within a 3D object is both more stealthy and more practical, while still preserving the object’s global geometric structure. These properties make patch-wise triggers a particularly attractive option in point cloud scenarios. Nonetheless, most existing patch-wise trigger designs were developed for 2D images, and their direct adaptation to 3D point clouds is non-trivial due to the unordered and irregular nature of point cloud data. This motivates our work, which explores the design of stealthy and optimized patch-wise triggers for 3D point-cloud models.
2.2 Backdoor Attacks and Defenses in 3D Point Clouds
Recent studies have explored backdoor attacks in 3D point clouds, proposing various trigger designs based on their unique properties. PointBA [16] introduced two attack strategies: PointBA-I, which injects an additional cluster of points as a trigger, and PointBA-O, which exploits a specific rotation angle as the backdoor trigger. Similar to PointBA-I, PCBA [37] embeds a spherical point cloud at an optimized location to manipulate model predictions. But the injected clusters are often visually conspicuous and can be effectively removed by augmentations. IRBA [7] improves stealth by generating sample-specific triggers through nonlinear local transformations, but it remains vulnerable to common data augmentations. IBAPC [5] optimizes a sample-wise spectral trigger, while iBA [2] requires training an extra folding-based autoencoder to generate geometric perturbations. Despite their different designs, both methods introduce significant computational cost that limits scalability and real-world applicability. SRBA [25] employs uniformly shifted input features as triggers, but this design alters the network’s input format, thereby limiting its generality. The above attack methods all rely on sample-wise trigger injection to generate poisoned samples, which inevitably introduces noticeable global geometric deformations or noise. In contrast, our approach is the first to optimize a local patch-wise trigger for 3D point cloud backdoor attacks, achieving favorable results in terms of attack effectiveness, stealthiness, and computational efficiency. Similar region-based perturbations have also appeared in recent adversarial attacks [39, 12]. Unlike these test-time, sample-specific methods, our approach is a train-time backdoor attack that embeds a unified patch-wise trigger transferable across samples. In addition, rather than relying on sample-level spectral analysis, we employ a curvature-based patch imperceptibility score for patch-level selection, ensuring efficiency and stealthiness.
On the defense side, conventional data augmentations are often employed as a defense strategy against backdoor attacks [7]. SOR [48] enhanced robustness by removing outliers from point clouds. Xiang et al. [38] proposed a reverse-engineering defense to detect the backdoored model. PointCRT [15] detected the poisoned samples by analyzing their abnormal corruption resistance. Lian et al. [21] introduced a novel cross-modal driven object restoration framework and mitigated the implanted triggers by reconstruction. Additionally, PointNCBW [35] explores watermarking in point clouds via a backdoor mechanism for positive purposes, which is beyond the scope of this work.
Figure 2: The framework of our proposed SPBA method. Given a benign point cloud sample $X$, SPBA first decomposes the sample into local patches and evaluates their patch imperceptibility scores (PIS). Patches with high PIS are then selected and transformed into the spectral domain using the Graph Fourier Transform (GFT). A spectral trigger is injected into the selected patches, ensuring stealthy perturbations. Finally, the Inverse Graph Fourier Transform (IGFT) reconstructs the local patches, which are then aggregated with the unmodified patches to form the final poisoned sample $X^{p}$.
3 Preliminaries
3.1 Problem Statement
We define the training dataset with $N$ samples as $D=\{(X_{i},Y_{i})\}_{i=1}^{N}$, where $P_{i}\in\mathbb{R}^{n\times 3}$ denotes the 3D coordinates of the $i^{th}$ sample, and $Y_{i}\in\{1,\dots,K\}$ is its corresponding label from a predefined set of $K$ classes. Building upon the previous explanation [10, 6], we divide the training dataset into two subsets: a benign subset $D_{b}$ and a poisoned subset $D_{p}$. The poisoned subset consists of a portion $\rho$ of the training dataset $D$, where selected samples are manipulated by the trigger injection function $\mathcal{T}$ (as shown in Fig. 2) and their labels are modified to a predefined target label $Y_{t}$. Given that the attacker has full control over the training process and access to the entire dataset, the backdoored model $f_{\theta}$ trained on this training dataset maintains high accuracy on benign samples while misclassifying any poisoned sample $X^{p}_{i}=\mathcal{T}(X_{i})$ as the target class $Y_{t}$.
3.2 Threat Model
Consistent with prior backdoor attack methods [37, 5], we assume that adversaries operate under a white-box setting, where they have full access to the model architecture, can manipulate a subset of the training data, and control the training process. By injecting poisoned samples into the training set, adversaries are able to implant backdoors into the DNNs. At the inference stage, we assume that adversaries are constrained to black-box interactions, i.e., they can only issue queries to the backdoored models with arbitrary inputs.
4 Method
4.1 Patchification and Curvature-Aware Selection
Due to the unordered and irregular nature of point clouds, it is infeasible to directly segment patches for trigger injection in the same manner as in 2D images. To address this challenge, we draw inspiration from previous works [44] and employ Farthest Point Sampling (FPS) and K-Nearest Neighbors (KNN) algorithms to systematically decompose each point cloud into local patches. This strategy enables a structured yet flexible partitioning of the point cloud into local patches. Each patch retains sufficient fine-grained geometric detail, supporting curvature-aware evaluation essential for spectral trigger embedding. Given an input point cloud sample $X$, FPS is first applied to select $g$ center points $X^{ct}=\{x^{ct}_{1},x^{ct}_{1},...,x^{ct}_{g}\}$, ensuring a well-distributed set of key reference points. For each center point $x^{ct}\in X^{ct}$, a local patch will be constructed by the KNN algorithm containing $k_{g}$ neighbors: $S_{i}=\mathit{K\!N\!N}(X,x_{i}^{ct})$ where $S_{i}\in\mathbb{R}^{k_{g}\times 3}$. The resulting set of patches is represented as $\{S_{1},S_{2},...,S_{g}\}$.
As human visual sensitivity varies across different regions of a 3D object, perturbations in high-complexity patches are less perceptible due to the absence of a fixed prior shape. We define a patch imperceptibility score (PIS), which draws from the imperceptibility score concept of the saliency and imperceptibility score (SI score) [23]. However, our PIS is specifically designed to focus on patch-wise geometric complexity without incorporating the saliency score. The local curvature around each point $x_{j}$ in a point cloud $X$ is defined as:
$C(x_{j};X)=\frac{1}{k_{c}}\sum_{q\in\mathcal{N}_{x_{j}}}|\langle\frac{q-x_{j}}{\|q-x_{j}\|_{2}},\mathbf{n_{x_{j}}}\rangle|,$ (1)
where $\mathcal{N}_{x_{j}}$ is defined as the set of $k_{c}$ nearest neighbors of point $x_{j}$ and $\mathbf{n_{x_{j}}}$ denotes its normal vector. The imperceptibility score (IS) of each point is further computed as the standard deviation of curvature between the point $x_{j}$ and its neighboring points $\mathcal{N}_{x_{j}}$ (with neighborhood size fixed as $k_{c}=10$, following prior work [23]), capturing local curvature fluctuations:
$IS(x_{j};X)=\sqrt{\frac{1}{k_{c}}\sum_{q\in\mathcal{N}_{x_{j}}}(C(q)-\frac{1}{k_{c}}\sum_{q\in\mathcal{N}_{x_{j}}}C(q))^{2}}.$ (2)
Using this measure, we construct an imperceptibility map for the entire point cloud. Finally, the patch imperceptibility score (PIS) of decomposed local patch $S_{i}$ is calculated by averaging the score of all points within the patch:
$\mathit{PIS}(S_{i})=\frac{1}{k_{g}}\sum_{x\in{S_{i}}}\mathit{IS}(x).$ (3)
From the decomposed patches, we select the patches $\{S^{\prime}_{1},S^{\prime}_{2},...,S^{\prime}_{m}\}$ with the top-$m$ ranked PIS for subsequent trigger injection, making perturbations imperceptible.
4.2 Patch-wise Spectral Trigger Injection
Designing local patch-wise triggers for 3D point clouds presents significant challenges due to their unordered and irregular structure, which limits the transferability of spatial local triggers across samples. When injected into different regions of different point cloud instances, these triggers often introduce noticeable outliers and cause severe semantic distortions, ultimately compromising both stealthiness and attack effectiveness.
To overcome these limitations, we propose designing local patch-wise triggers in the spectral domain [14, 22]. Specially, a selected point cloud patch $S^{\prime}\in\{S^{\prime}_{1},S^{\prime}_{2},...,S^{\prime}_{m}\}$ is first transformed into an unweighted KNN graph $G=(V,A)$. Here $V\in\mathbb{R}^{k_{g}\times 3}$ represents the set of vertices, where each vertex corresponds to a point in the patch $S^{\prime}$. Following prior works [5, 14, 22], the adjacency matrix $A\in\mathbb{R}^{k_{g}\times k_{g}}$ is constructed by connecting each point to its $k_{p}=10$ nearest neighbors for graph construction, where $A_{ij}=1$ if points $i$ and $j$ are connected, otherwise $A_{ij}=0$.
From this, we compute the combinatorial graph Laplacian matrix [31] defined as L=DAL=D-A, where $D\in\mathbb{R}^{k_{g}\times k_{g}}$ is the degree matrix, with each entry defined as $D_{ij}=\sum_{j=1}^{N}A_{ij}$. Since $L$ is real, symmetric, and positive semi-definite, we perform eigen decomposition on it as $L=U\Lambda U^{T}$, where $U\in\mathbb{R}^{k_{g}\times k_{g}}$ is an orthonormal matrix containing the eigenvectors and $\Lambda$ consists of eigenvalues.
Using this decomposition, we compute the spectral feature of $S^{\prime}$ via the Graph Fourier Transform (GFT), where low-spectrum components capture coarse structural information, while high-spectrum components encode fine-grained geometric details:
$\hat{S^{\prime}}=\phi_{GFT}(S^{\prime})=U^{T}S^{\prime},$ (4)
where $\hat{S^{\prime}}\in\mathbb{R}^{k_{g}\times 3}$. Correspondingly, the Inverse GFT (IGFT) could reconstruct the spatial feature from the spectral domain:
$S^{\prime}=\phi_{IGFT}(\hat{S})=U\hat{S^{\prime}}.$ (5)
Then, we propose an optimizable spectral trigger $\xi\in\mathbb{R}^{k_{g}\times 3}$, which is applied to the spectral features of selected patches to generate the poisoned patches $S^{p}$:
$S^{p}=\phi_{IGFT}(\phi_{GFT}(S^{\prime})+\xi).$ (6)
The final poisoned point cloud sample $X^{p}$ is then constructed by replacing the original selected patches with the reconstructed patches while keeping the unselected regions unchanged:
$X^{p}=\mathcal{T}(X)=(X\setminus\{S^{\prime}_{1},S^{\prime}_{2},...,S^{\prime}_{m}\})\cup\{S_{1}^{p},S_{2}^{p},...,S_{m}^{p}\}.$ (7)
Figure 3: Comparison of original and perturbed patches of airplane tail. The second row shows the spectral features of the original patch and spectral perturbations across different spectrum bands.
4.3 Optimization Strategy
As illustrated in Fig. 3, low-spectrum perturbations cause significant deformation to the patch of empennage, while high-spectrum perturbations introduce excessive outliers. Perturbations applied within intermediate spectral bands mitigate outliers while preserving the overall shape, presenting a key challenge for the optimal design of spectral triggers. To address this, we propose an optimizable spectral trigger $\xi$ that strikes a balance between attack effectiveness and stealthiness. The backdoor learning process follows an alternating optimization scheme, where the model parameters $\theta$ and the spectral trigger $\xi$ are updated iteratively, with one being optimized while the other remains frozen. Thus, each optimization step consists of two forward and backward propagations.
To train the model parameters $\theta$, we define the loss function $L_{M}$, following the standard formulation for backdoor attacks:
$L_{M}=\sum_{D_{b}}L_{CE}(f_{\theta}(X),Y)+\sum_{D_{p}}L_{CE}(f_{\theta}(X^{p}),Y_{t}),$ (8)
where $L_{CE}$ represents the cross-entropy loss function.
To optimize the patch-wise trigger $\xi$ for both attack efficiency and high stealthiness, we ensure that poisoned samples are misclassified into the target class $Y_{t}$ while maintaining minimal perceptual distortions. This is achieved through the following loss function $L_{T}$:
$L_{T}=\sum_{D}\left(L_{CE}(f_{\theta}(\mathcal{T}(X)),Y_{t})+L_{reg}(X,\mathcal{T}(X))\right).$ (9)
In the regularizer function $L_{reg}$, we employ Euclidean distance $L_{ED}(X,\mathcal{T}(X))$ and Chamfer distance $L_{CD}(X,\mathcal{T}(X))$ to enforce similarity between the original and poisoned point clouds, and Hausdorff distance $L_{HD}(X,\mathcal{T}(X))$ is used to control the worst scenario of outliers. The final regularization term is formulated as:
$L_{reg}=\lambda_{1}L_{ED}+\lambda_{2}L_{CD}+\lambda_{3}L_{HD},$ (10)
where $\lambda_{1},\lambda_{2},\lambda_{3}$ control the relative contribution of each loss term. This joint optimization framework ensures that the spectral trigger remains both effective and imperceptible, enabling stealthy backdoor attacks on 3D point cloud models.
5 Experiments and Results
5.1 Experiment Details
Datasets and Models. We evaluate our proposed method on two widely used 3D point cloud classification datasets: 1) ModelNet40 [36]: a benchmark dataset containing 40 object categories, with an official split of 9,843 training samples and 2,468 testing samples; 2) ShapeNetPart [41]: a subset of ShapeNet, consisting of 12,128 training samples and 2,874 testing samples across 16 object categories. For both datasets, we uniformly sample 1,024 points from each shape along with their normal vectors, normalizing the coordinates for consistency. For backdoored models, we adopted the most widely used models in previous work [16, 7], including PointNet [26], PointNet++ [27], DGCNN [34], and PointNext [28].
Attacks Setup. We compare our proposed SPBA method against four representative point cloud attack methods: PointBA-I [16], PointBA-O [16], IRBA [7], and IBAPC [5]. To reproduce PointBA-I, we implant a spherical cluster comprising 3% of the total points, with a fixed radius of 0.05, positioned at $(0.5,0.5,0.5)$ in the coordinate space. For PointBA-O, we apply a $10^{\circ}$ rotation along the Z-axis. For IRBA and IBAPC, we follow the default configurations specified in their original papers. For our proposed SPBA, $g$ and $k_{g}$ are both set as 32 for patch decomposition. And $m=16$ patches are selected for trigger injection. For the regularizer function in Eq. (10), we set $\lambda_{1}=1$, $\lambda_{2}=5$ and $\lambda_{3}=1$. However, for PointNet, we adjust $\lambda_{1}$ to 0.1 to improve stability.
For all attack methods, we set the poisoning rate $\rho=0.1$. The target class is chosen as Chair ($Y_{t}=8$) for ModelNe40 and Lamp ($Y_{t}=8$) for ShapeNetPart. We employ two Adam optimizers for backdoored model training and spectral trigger optimization, with learning rates of 0.001 and 0.01, respectively, and a weight decay of 1e-4. A cosine annealing learning rate scheduler is applied to both optimizers to facilitate stable convergence. Training is performed for 100 epochs with a batch size of 32. For all experiments, we use a workstation equipped with an NVIDIA RTX 4090 GPU.
Evaluation Metrics. The attack performance can be evaluated by Benign Accuracy (BA) and Attack Success Rate (ASR). BA is the accuracy of benign test samples that are correctly classified. ASR represents the proportion of poisoned test samples that are successfully misclassified into the target label. To quantify the imperceptibility of poisoned samples, we compute the Chamfer Distance (CD) between the poisoned and original samples. A successful backdoor attack should achieve a high ASR and lower CD while preserving BA as much as possible. For clarity, both BA and ASR are presented as percentages, and CD values in tables are scaled by a factor of 1,000.
Table 1: Comparison of attack performance across different models on ModelNet40. BA and ASR are presented as percentages, and CD values in tables are scaled by a factor of 1,000. The best results are bolded and the second-best results are underlined.
Model Clean PointBA-I PointBA-O IRBA IBAPC SPBA
[…]
PointNet 89.14 89.67 99.03 4.08 88.17 94.77 5.31 88.25 93.27 18.54 88.21 91.57 3.04 88.90 97.89 2.60
PointNet++ 90.80 90.36 98.87 4.08 89.55 96.68 5.31 90.40 98.26 18.54 89.67 95.58 0.68 89.42 97.97 0.11
DGCNN 89.75 91.05 100.00 4.08 89.67 95.02 5.31 89.55 92.38 18.54 88.24 95.44 0.87 89.75 97.61 0.69
PointNext 89.22 88.21 95.83 4.08 88.94 96.31 5.31 88.41 92.46 18.54 88.70 96.60 1.48 88.62 96.85 0.92
Average 89.73 89.82 98.43 4.08 89.08 95.69 5.31 89.15 94.09 18.54 88.70 94.80 1.52 89.17 97.58 1.08
5.2 Attack Effectiveness
The results presented in Table 1 provide a comparative evaluation of various backdoor attack methods on ModelNet40. As shown in Table 1, our proposed SPBA achieves the lowest average CD value of 1.08 among all attacks, demonstrating its state-of-the-art stealthiness. What’s more, SPBA outperforms most backdoor attacks according to BA and ASR. In particular, compared to the sample-wise spectral attack IBAPC, SPBA achieves notable improvements of 0.47% in average BA and 2.78% in average ASR, highlighting the effectiveness of our patch-wise spectral trigger. Although PointBA-I achieves the highest average ASR of 98.43%, it comes at the cost of a significantly higher average CD of 4.08, making it much more detectable. Subsequent defense experiments further confirm that the trigger used in PointBA-I is especially susceptible to detection or removal by existing defense methods. Table 2 shows more experimental results on the ShapeNetpart dataset. SPBA achieves a similar performance, offering the best balance between attack effectiveness and stealthiness. These results further validate the superiority of SPBA as a more effective and imperceptible backdoor attack method compared to existing attacks.
Table 2: Comparison of attack performance across different models on ShapeNetPart. BA and ASR are presented as percentages, and CD values in tables are scaled by a factor of 1,000. The best results are bolded and the second-best results are underlined.
Model Clean PointBA-I PointBA-O IRBA IBAPC SPBA
[…]
PointNet 98.47 98.30 99.72 7.48 97.95 98.34 5.67 98.12 96.80 16.06 97.80 96.01 1.73 98.30 98.40 1.22
PointNet++ 98.99 98.75 99.69 7.48 98.50 98.78 5.67 98.89 99.03 16.06 98.43 98.46 0.22 98.26 98.30 0.02
DGCNN 99.06 98.92 100.00 7.48 98.85 98.50 5.67 98.68 97.63 16.06 97.54 95.44 0.80 98.43 99.23 0.44
PointNext 98.82 98.61 99.51 7.48 98.10 96.83 5.67 96.31 96.17 16.06 97.35 95.99 0.44 98.53 98.45 0.37
Average 98.83 98.64 99.73 7.48 98.35 98.11 5.67 98.00 97.41 16.06 97.78 96.48 0.80 98.38 98.60 0.51
5.3 Attack Stealthiness
We visualize some poisoned samples generated by different backdoor attack methods in Fig. 4. As observed, PointBA-I introduces noticeable anomalies compared to clean samples, significantly increasing its detectability. Although PointBA-O and IRBA generate natural-looking poisoned samples by applying transformation, they still cause considerable deformations in shape. Compared to IBAPC, SPBA selectively perturbs the flower region while preserving the smooth structure of the vase, making the modifications less perceptible to human observers.
Furthermore, we conduct a human perceptual study to assess the stealthiness of the proposed attack. Following the protocol in IRBA [7], we present a mixture of clean and poisoned samples to 20 participants, all of whom are over 18 years old, major in computer science, and have prior knowledge of machine learning and computer vision. Each participant is shown 10 groups of multi-view renderings of point clouds from distinct object classes. Each group consists of one clean reference sample together with five attack types of its poisoned version arranged in random order. Participants are instructed to assign a score $\in\{1,2,3,4,5\}$ to indicate the likelihood that the poisoned samples in the group have been modified. A lower score reflects a lower perceived likelihood, suggesting that the poisoned samples are less perceptible. The average score for each attack method is reported in Table 3. The results demonstrate that SPBA achieves the highest level of stealthiness, yielding the lowest perceptibility score among all compared methods.
Figure 4: Visual comparison of poisoned samples from different backdoor attack methods. Our proposed SPBA preserves structural integrity and ensures smooth regions (e.g., the vase or car roof) remain undisturbed, enhancing imperceptibility.
Table 3: Scores of the human perceptual study.
Method PointBA-I PointBA-O IRBA IBAPC SPBA
Score $\downarrow$ 4.79 3.33 4.05 2.76 1.53
5.4 Attack Efficiency
As shown in the Eq. (6), the computational cost of spectral attacks is highly dependent on the trigger size. And we further report the computational cost of different spectral attacks (with a batch size of 64 during the inference stage) in Table 4. The prior spectral attack IBAPC uses a large trigger of size $\xi^{\text{IBAPC}}\in\mathbb{R}^{1024\times 3}$, resulting in a substantial computational cost of 939.72 MFLOPs, which severely hinders its efficiency and practicality for deployment. In contrast, although our proposed SPBA perturbs an average of 452 points, which account for 44% of the sample, our patch-based spectral trigger is much smaller ($\xi^{SPBA}\in\mathbb{R}^{32\times 3}$). The computational cost of SPBA is only 14.74 MFLOPs, which is comparable to that of a single fully connected layer. Furthermore, similar to IBAPC, we construct a sample-based spectral trigger using the FPS algorithm (denoted as FPSP, $\in\mathbb{R}^{452\times 3}$) that perturbs the same number of points as SPBA. However, it significantly increases the computational cost by a factor of 11.54 compared to our patch-based trigger. Moreover, it results in weaker attack performance (shown in Table 8), highlighting the advantage of our patch-wise design. These results demonstrate the efficiency and effectiveness of our proposed patch-wise spectral trigger.
Table 4: Computational cost of spectral attacks.
Method IBAPC FPSP SPBA
MFLOPS $\downarrow$ 939.72 170.06 14.74
5.5 Resistance to Defense Methods
Resistance to Data Augmentation. Since data augmentation is widely employed to enhance the resistance of classification models, we introduce six types of augmentations to evaluate their impact on backdoor attack performance: 1) Rotation: random rotation around the z-axis with a maximum angle of $10^{\circ}$; 2) Rotation3D: random rotation up to $10^{\circ}$ along all three axes; 3) Scaling: scaling within a range of 0.5 to 1.5; 4) Shift: shifting randomly between -0.1 and 0.1 along each axis; 5) Dropout: removing up to 20% of points; and 6) Jitter: introducing noise sampled from $N(0,0.02)$. The experiments are conducted on the ModelNet40 dataset using the PointNet++ model. As shown in Table 5, different augmentations affect attack methods to varying degrees. Due to the high visibility of the trigger of PointBA-I, augmentations have minimal impact, as the trigger remains distinct. However, for rotation-based PointBA-O, applying Rotation significantly reduces attack effectiveness. For IRBA, which relies on local transformations, augmentations such as Rotation3D, Dropout and Jitter cause a notable decline in attack performance, with ASR dropping by up to 8.23%. Conversely, for optimization-based attacks like IBAPC and our proposed SPBA, only Jitter leads to a slight ASR reduction, while other augmentations have negligible effects, demonstrating their resistance against most augmentations.
Table 5: Resistance of backdoor attacks to varied data augmentations and SOR.
Augmentation PointBA-I PointBA-O IRBA IBAPC SPBA
[…]
No 90.36 98.87 89.55 96.68 90.40 98.26 89.67 95.58 89.42 97.97
Rotation 89.42 98.95 86.83 7.66 90.52 97.69 90.59 95.90 89.34 98.22
Rotation3D 89.71 99.11 88.78 30.51 90.68 92.79 90.23 95.94 89.38 97.24
Scaling 90.72 99.55 91.65 96.92 91.17 99.47 90.85 96.47 91.57 98.26
Shift 90.11 99.03 89.67 96.07 90.03 97.93 89.74 94.23 89.59 97.08
Dropout 90.52 99.39 90.07 96.11 90.40 94.85 88.37 94.93 90.76 97.24
Jitter 89.42 98.70 89.34 95.06 90.07 90.03 89.91 91.65 90.96 95.01
SOR 90.36 46.47 89.55 95.46 90.40 97.30 89.67 95.13 89.42 97.36
Resistance to SOR. Statistical Outlier Removal (SOR) [48] is designed to eliminate outliers based on the local neighborhood density using the KNN algorithm. In our implementation, we set KNN to use 60 nearest neighbors and set the standard deviation threshold to 0.5. We apply it during inference on PointNet++ models trained on ModelNet40. The results in Table 5 show that only PointBA-I experiences a significant drop in ASR, decreasing from 98.87% to 36.50%, while the ASR reduction for other attacks remains within 2%. This suggests that SOR serves as an effective defense specifically against injection-based attacks while having a limited impact on other attack methods.
Resistance to the Gradient-based Method. As many studies on backdoor attacks assess trigger stealthiness using gradient-based methods [30, 24], we employ a point cloud saliency map [47] to identify salient points that contribute most to the model’s decision-making. Specifically, 40 points with the highest saliency scores are marked in red. As illustrated in Fig. 5, the spherical trigger of PointBA-I is easily detected. In contrast, our proposed SPBA exhibits a saliency distribution similar to the clean model, indicating enhanced stealthiness and reduced detectability.
Figure 5: Gradient-based salience analysis with the most significant points highlighted in red.
Resistance to PointCRT. PointCRT [15] utilizes a nonlinear classifier to detect the abnormal corruption resistance of poisoned samples. The F1 score (F1) and Area Under the Curve (AUC) are employed to evaluate the ability of PointCRT to detect the attack, where lower values indicate that the attack is more resistant to detection by PointCRT. As shown in Table 6, SPBA outperforms almost all the attacks except PointBA-O, which exhibits limited attack performance and is vulnerable to data augmentations.
Table 6: Performance of PoinCRT on varies attacks.
Method PointBA-I PointBA-O IRBA IBAPC SPBA
F1 $\downarrow$ 95.94 72.89 93.98 92.84 87.76
AUC $\downarrow$ 97.52 77.76 95.75 92.93 87.89
Resistance to FLARE. FLARE [13] aggregates abnormal activations across hidden layers and adaptively selects the most stable subspace to divide samples into two clusters, detecting the more stable cluster as poisoned samples. Since this method is independent of input modality, it could also be applied to point cloud backdoor sample detection. Specifically, we evaluated the method on backdoored DGCNN models trained on ModelNet40. The true positive rate (TPR) reflects the proportion of correctly identified poisoned training samples, with a lower TPR indicating weaker detection capability. Experimental results in Table 7 show that our approach consistently yields the lowest TPR, demonstrating strong resistance against FLARE.
Table 7: Detect accuracy of FLARE on varies attacks.
Method PointBA-I PointBA-O IRBA IBAPC SPBA
TPR $\downarrow$ 56.93 62.01 47.97 38.01 29.78
5.6 Ablation Study
Effect of Patch Selection Strategies. In our default setting, we select the 16 patches with the highest patch imperceptibility scores (PIS), denoted as H-PSI. To evaluate the effectiveness of this selection strategy, we compare H-PSI with three alternative patch selection strategies: (1) L-PSI: Select the patches with the lowest PIS; (2) Random: Randomly select the patches; (3) FPSP: Use the FPS algorithm to sample the same number of points as H-PSI. We evaluate their average performance on ModelNet40. As shown in Table 8, the H-PSI strategy achieves the highest attack effectiveness and the strongest stealthiness. This result underscores the advantage of selecting high-imperceptibility patches for local spectral trigger injection.
Table 8: Comparison of attack performance of SPBA with various selection strategies.
Method L-PSI Random FPSP H-PSI(default)
BA $\uparrow$ 87.99 88.19 87.84 89.17
ASR $\uparrow$ 95.53 96.35 96.45 97.58
CD $\downarrow$ 1.16 1.29 1.39 1.08
Effect of Selected Patch Number $m$. In this section, we experiment with injecting the spectral trigger into varying numbers $m$ of selected patches during the training process. The experiments are conducted on ModelNet40 using PointNet++. As shown in Table 9, as $m$ increases from 1 to 16, the CD value gradually decreases, indicating improved stealthiness. This suggests that selecting more patches within a reasonable range helps distribute perturbations more naturally, enhancing imperceptibility. However, when $m$ further increases to 32, where almost all points are perturbed, CD rises to 0.12, indicating that excessive modifications start to introduce noticeable distortions, reducing stealthiness. Based on these observations, we set $m=16$ as the default configuration, as it achieves the optimal balance between imperceptibility (lowest CD), strong attack effectiveness, and benign accuracy.
Table 9: The effect of selected patch number $m$.
$m$ 1 2 4 8 16 24 32
BA 89.56 89.88 88.53 90.15 89.42 89.27 89.36
ASR 97.38 97.10 97.28 96.84 97.97 97.73 97.30
CD 0.17 0.19 0.18 0.14 0.11 0.11 0.12
Effect of the Patch Size. We conduct additional experiments on ModelNet40 using PointNet++ to further investigate the effect of patch size. To ensure that the total number of perturbed points remains approximately constant, when the patch size is doubled, the number of selected patches is reduced to half, and vice versa. From Table 10, ASR attains its maximum at $k_{g}=32$ and CD is minimized at this setting, offering the best balance of attack success and stealth. The benign accuracy is still high (89.42%), ranking second among the evaluated patch sizes.
Table 10: The effect of patch size $k_{g}$
$k_{g}$ 16 32 64 128
BA 89.30 89.42 89.75 89.26
ASR 97.00 97.97 97.64 97.73
CD 0.13 0.11 0.18 0.15
Effect of the Target Class. In this section, we investigate the effect of different target classes on our attack performance. Specifically, we select the top 8 categories from ModelNet40 as target classes and conduct experiments using PointNet++. As shown in Fig. 6(a), the choice of target class has a minimal effect on both Attack Success Rate (ASR) and Benign Accuracy (BA), with fluctuations not exceeding 1%. These results demonstrate the robustness of our approach across various target label settings, ensuring its consistency in different attack scenarios.
(a) The effect of the target class selection.
(b) The effect of poisoning rate.
Figure 6: The effect of the target class selection and poisoning rate.
Effect of the Poisoning Rate. We explore the impact of the poisoning rate on our attack by conducting experiments on ModelNet40 using PointNet++. As shown in Fig. 6(b), ASR consistently increases as the poisoning ratio rises. BA exhibits minor fluctuations but remains relatively stable as the poisoning rate increases. Notably, even at a low poisoning rate of 2.5%, our approach achieves an ASR of 96.27%, underscoring its effectiveness even with limited poisoned data.
Effect of the Regulation Loss Weights. In this section, we investigate the effect of the regulation loss weights in Eq. 10, where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ control the contributions of $L_{ED}$, $L_{CD}$, and $L_{HD}$, respectively. As shown in Table 11, increasing $\lambda_{1}$ gradually enhances the stealthiness (reflected by lower CD values), but at the cost of a decreasing ASR. Notably, when $\lambda_{1}=1$, the attack achieves a relatively good trade-off between aggressiveness and stealthiness. Moreover, our proposed SPBA algorithm demonstrates robustness with respect to the hyperparameters $\lambda_{2}$ and $\lambda_{3}$. Based on these observations, we ultimately set $\lambda_{2}=5$ and $\lambda_{3}=1$ to achieve a favorable balance between attack effectiveness and stealthiness.
Table 11: The effect of the regulation loss weights
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ BA ASR CD
1 5 1 89.42 97.97 0.11
0.1 5 1 89.72 99.39 0.72
0.5 5 1 89.47 98.50 0.19
2 5 1 87.64 92.70 0.08
1 1 1 89.95 97.17 0.15
1 3 1 89.34 97.77 0.11
1 5 0.1 89.74 96.96 0.12
1 5 0.5 89.62 97.44 0.11
1 5 2 89.52 97.71 0.11
6 Conclusion
In this paper, we propose SPBA, a novel backdoor attack method against 3D point cloud models. Unlike existing methods that rely on sample-wise trigger injection, which introduces noticeable geometric distortions, SPBA adopts a local patch-wise spectral trigger that selectively perturbs geometrically complex regions, demonstrating superior stealthiness, efficiency and effectiveness. Additional experiments validate the strong resistance of SPBA to existing defense methods. Experimental results highlight the effectiveness of patch-wise triggers as a promising direction for stealthy and adaptive backdoor attacks in 3D point clouds.